Please redo whole problem

5)You are a consultant to a mid-sized manufacturing corporation that is considering an investment project. The project requires an initial investment of $100 million and will generate an after tax cash of $20 million in the first year and the cash flow will increase 5% thereafter every year (Please note that this is a constant growing cash flow).The project’s beta is 1.5. Assuming that rf=5% and E ( rM ) = 12%, Please answer the following questions.
What is the net present value of the project ?
What is the highest possible discount rate for the project before its NPV becomes negative ?
What is the highest possible beta estimate for the project before its NPV becomes negative ?
Answer #1

Calculating Required Rate using CAPM Model,

Required Rate = 0.05 + 1.5(0.12 - 0.05) = 15.50%

NPV = -100 + 20/(0.155 - 0.05)

NPV = $90.48 million

Calculating Highest possible discount rate,

100 = 20/(r - 0.05)

r = 25.0%

Calculating beta,

0.25 = 0.05 + Beta(0.12 - 0.05)

Beta = 2.86